OpenStudy (anonymous):

Wilson invests $8000 in an account that pays 6.4% per year compounded quarterly. a) Find the amount in the account after 3 years. b) How long will it take for the investment to double? c) how long will it take for the amount to be $10500? I don't really know how to do this.. please help! thanks a lot!!

OpenStudy (anonymous):

do you know the formula for compound interest?

OpenStudy (anonymous):

um... A=P(1+r/n)^nt??

OpenStudy (anonymous):

isn't it?

OpenStudy (anonymous):

\[Compound interest = principle * ( 1 + ((rate/100)^t))\]

OpenStudy (amistre64):

Keita, you are right :)

OpenStudy (anonymous):

keita you are right. so what is n in your equation, what is P , what is r and what is t?

OpenStudy (anonymous):

So it is 8000 * ( 1.064 ^ 12) = 16841.84

OpenStudy (anonymous):

For the first part

OpenStudy (amistre64):

8000 (1 + .064/4)^(4*t)

OpenStudy (amistre64):

ishan; compunded quarterly means that the yearly interest rate is divided up and paid out 4 times a year right?

OpenStudy (anonymous):

yup my bad i missed the quarterly

OpenStudy (amistre64):

:) i do that all the time ;)

OpenStudy (anonymous):

Using amistres formula it is 8000 ( 1 + 0.64/4) * 4 ^ t = 16000 4^t * 1.016 * 8000 = 4^t * 8128 = 16000 4^t = 7872 t = log 7872 to the base 4 t = 6.47 years or 6 years 5 months and 20 days

OpenStudy (anonymous):

um... so the amount after 3 years is.. 16000 or 7872?? and how long will it take to be double?